OpenStudy (abbycross167):

Can someone please help me fix this math question? Please? I will fan and give a medal!! Please help me to simplify this rational expression - ax+bx 2x ------ / ------ x2 a+b **Note the / means to divide

OpenStudy (abbycross167):

these are the steps I did but my teacher said they are incorrect -- Steps- ax+bx 2x ------ / ------ x2 a+b x(a+b)*2x ---------- x2(a+b) = x*2x ----- x2 = 2

OpenStudy (blurbendy):

let's start with the top first. ((ax + bx) / (x^2)) this can be rewritten as ax / x^2 + bx / x^2 do you follow so far?

OpenStudy (abbycross167):

yes sir

OpenStudy (blurbendy):

okay, first can you simplify ax / x^2 ?

OpenStudy (abbycross167):

I don't think so

OpenStudy (blurbendy):

you can since there is an x in the denominator and numerator

OpenStudy (blurbendy):

so, what does the simplified expression become?

OpenStudy (abbycross167):

oh ok

OpenStudy (abbycross167):

would it be a^2?

OpenStudy (blurbendy):